OpenStudy (anonymous):

I know this is elementary type stuff but i just need a reminder. Convert 0.5% to a decimal.

OpenStudy (amistre64):

1% = 1/100 mutliply both sides by .5

OpenStudy (anonymous):

@amistre64 im supposed to move the decimal to the right or left ?

OpenStudy (anonymous):

i move the decimal to the left correct? @amistre64

OpenStudy (anonymous):

When converting anything to a decimal from percentages, move the decimal two places to the left

OpenStudy (anonymous):

lol so what i said haha.

OpenStudy (anonymous):

And if you are given a decimal and asked to convert it into a percentage, move it to the right two places

OpenStudy (anonymous):

right thank you @SymbioteBowTie i just needed a little reminder thx

OpenStudy (anonymous):

@SymbioteBowTie how would i convert 15 over 50 to a percent?

OpenStudy (anonymous):

15/50?

OpenStudy (anonymous):

I personally, if this is a fraction, would take 15/50 and 15 ÷ 50 to get .3, then move the decimal to places to the right. Then I'd be left with 30%

OpenStudy (anonymous):

15 divided by 50* Sorry, the questionmark symbols are annoying...

OpenStudy (anonymous):

ohh wow so simple but i just cant dont remember the formulas :/ lol @SymbioteBowTie

OpenStudy (anonymous):

It's alright, one more tip:

OpenStudy (anonymous):

You can easily check this by taking .3 and multiplying it with 50 and you would get 15

OpenStudy (anonymous):

This way you can tell that 15 is 30% of 50

OpenStudy (anonymous):

ok i have 3/8 so is 0.375 so then i move the decimal to the right 2 times and i have 37.5% correct?

OpenStudy (anonymous):

@SymbioteBowTie

OpenStudy (anonymous):

@SymbioteBowTie and can you tell me how is/of=%/100 formula works

OpenStudy (amistre64):

moving decimals is not a math technique, it relies on you to memorize some rule instead of relying on some rather basic math skills to accomplish the same task.

OpenStudy (amistre64):

1% = 1/100 = .01 .5(1%) = .5(.01) = .005